A grocer mixes peanuts that cost $1.49 per pound and walnuts that cost $2.69 per pound to make 100 pounds of a mixture that cost $2.21 per pound.How much of each kind of nut is put into the mixture?
Answer by richwmiller(17219)

total sold 100
total sales 221
product sold at 1.49 and 2.69
p +w =100,
1.49*p +2.69*w =221
p =100-w
We have two equations with two variables.
We will substitute to make one equation with one variable
1.49*(100-w )+2.69*w =221
149.0-1.49w +2.69*w =221
1.2*w =72.0
w =60.0
p =100-w
p =40 at $1.49 and w =60 at $2.69
check
40+60=100
1.49*p +2.69*w =221
1.49*40+2.69*60=221
59.6+161.4=221
221.0=221
ok
codetickv